calculate the final pH when 120.0mL of 0.50M HI is added to 850 mL of 0.05M HClO4
1
Expert's answer
2020-05-03T15:22:58-0400

Solution.

"V(HI) = 120.0mL=0.120L;"

"M(HI) =0.50M;"

"V(HClO_4) =850mL=0.850L;"

"M(HClO_4)=0.05M;"

"\nu(HI)=V(HI)\sdot M(HI); \nu(HI)=0.120L\sdot0.50M=0.06mol;"

"\nu(HClO_4)=V(HClO_4)\sdot M(HClO_4);"

"\nu(HClO_4)=0.850L\sdot0.05M=0.0425mol;"

"\nu_{final}=0.06mol+0.0425mol=0.1025mol;"

"[H_3O^+]=\dfrac{\nu_{final}}{V_{final}}; [H_3O^+]=\dfrac{0.1025mol}{0.970L}=0.1057M;"

"pH=-lg[H_3O^+]=-lg(0.1057)=lg9.46=0.976" ;

Answer: "pH=0.976."